# Sylvio Ferraz-Mello 11    Cristian Beaugé 22

### v1

@@ 3 Orbital evolution of a planet initially in the 3:2 spin-orbit resonance
Figures 4 and 5 show rapid damping of the eccentricity and escape from the 3:2 spin-orbit resonance. We see two different phases in this process. First, the eccentricity is quickly damped to a small value. When the decreasing eccentricity reaches the critical value corresponding to the spin-orbit resonance (see Appendix A.1), the stable stationary solution ceases to exist, and the rotation escapes the resonance and is slowly driven to synchronization. This phase lasts a few thousand years.
We note that in this phase, the orbital period is increasing. That is, $\dot{a}>0$. We recall that the contribution of the planetary tide to the variation of the semi-major axis when the planet’s rotation is faster than the orbital motion is positive and, in the first approximation, independent of the eccentricity:
@@
where $\nu_{p}=2(\Omega_{p}-n)$ is the so-called semi-diurnal frequency. This equation is not enough to understand the cases in the 3:2 spin-orbit resonance, but it applies to what is seen in the first thousand years of Fig. 5 because the planet escapes the resonance and the eccentricity has a fast decrease. Otherwise, a more complete expression is necessary (see Appendix B).
@@ 4 Heat Dissipation
The variations may be determined from the quantities $\dot{P}$ obtained from the results shown in Figs. 5 and 6. As discussed in previous sections, the planetary tide is dominant when the eccentricity is forced to remain constant, and so the planet is the body responsible for the orbital energy loss.
@@ Appendix B Variation of the semi-major axis
The first three terms in brackets do not present difficulties. In these cases, we have $\gamma_{p}\ll|X|$ and may write $\Psi(X)\simeq\frac{\gamma_{p}}{X}$. However, the fourth term needs a more detailed calculation. If the eccentricity is large enough to allow the crest formed near the point $\Omega_{p}=\frac{3n}{2}$ (see Fig. 12) to cross the axis $\dot{\Omega}_{p}=0$, the equation $\dot{\Omega}_{p}=0$ has two real roots. They are $\nu_{p}-n\simeq-\frac{49}{4}ne^{2}$ (unstable) and $\nu_{p}-n\simeq-\frac{4\gamma_{p}^{2}}{49ne^{2}}$ (stable). If $\gamma_{p}\ll\frac{49}{4}ne^{2}$, we get the approximate result $\frac{147e^{2}}{8}\Psi(\nu_{p}-n)\simeq-\frac{3}{2}$, which makes $\dot{a}_{p}<0$. Clearly, this result is not true for a very small eccentricity, but it is valid in the case under study and explains the behavior observed by the evolution curves in Fig. 6. It is not visible in Fig. 5, notwithstanding the fact that the initial rotation is in the 3:2 spin-orbit resonance, because the escape of this condition is very fast and not included in the timespan of the figure. When the solution reaches the lower value on the shown time range ($t=1$ kyr), it is no longer trapped in the resonance.
@@ Appendix C Used parameters
Table 1: Other stellar and planetary parameters adopted
@@
Relaxation factor (s-1) $50-200$ $0.2-7\times 10^{-7}$
Moment of Inertia (×$\times\ mR^{2}$) 0.07 0.33
@@
Rotation Period (days) 42.7 --

### v2

@@ 3 Orbital evolution of a planet initially in the 3:2 spin-orbit resonance
Figures 4 and 5111Figure 5 replaces the one originally published, which was flawed due to a coding mistake. The text discussing the figure has been changed accordingly. show rapid damping of the eccentricity and escape from the 3:2 spin-orbit resonance. We see two different phases in this process. First, the eccentricity is quickly damped to a small value. When the decreasing eccentricity reaches the critical value corresponding to the spin-orbit resonance (see Appendix A.1), that stable stationary solution ceases to exist, and the rotation escapes the resonance and is driven to synchronization. While trapped in the resonance, $\dot{a}>0$.
We recall that the contribution of the planetary tide to the variation of the semi-major axis when the planet’s rotation is faster than the orbital motion is positive and, in the first approximation, independent of the eccentricity:
@@
where $\nu_{p}=2(\Omega_{p}-n)$ is the so-called semi-diurnal frequency. This equation is not enough to understand the cases in the 3:2 spin-orbit resonance, but it applies to what is seen in Fig. 5 because the planet escapes the resonance.
@@ 4 Heat Dissipation
The variations may be determined from the quantities $\dot{P}$ obtained from the results shown in Fig. 6. As discussed in previous sections, the planetary tide is dominant when the eccentricity is forced to remain constant, and so the planet is the body responsible for the orbital energy loss.
@@ Appendix B Variation of the semi-major axis
The first three terms in brackets do not present difficulties. In these cases, we have $\gamma_{p}\ll|X|$ and may write $\Psi(X)\simeq\frac{\gamma_{p}}{X}$. However, the fourth term needs a more detailed calculation. If the eccentricity is large enough to allow the crest formed near the point $\Omega_{p}=\frac{3n}{2}$ (see Fig. 12) to cross the axis $\dot{\Omega}_{p}=0$, the equation $\dot{\Omega}_{p}=0$ has two real roots. They are $\nu_{p}-n\simeq-\frac{49}{4}ne^{2}$ (unstable) and $\nu_{p}-n\simeq-\frac{4\gamma_{p}^{2}}{49ne^{2}}$ (stable). If $\gamma_{p}\ll\frac{49}{4}ne^{2}$, we get the approximate result $\frac{147e^{2}}{8}\Psi(\nu_{p}-n)\simeq-\frac{3}{2}$, which makes $\dot{a}_{p}<0$. Clearly, this result is not true for a very small eccentricity, but it is valid in the case under study and explains the behavior observed by the evolution curves in Fig. 6. It is not visible in Fig. 5, notwithstanding the fact that the initial rotation is in the 3:2 spin-orbit resonance, because the escape of this condition is very fast.
@@ Appendix C Used parameters
Table 1: Stellar and planetary parameters adopted
@@
Mass 0.9 M 0.02703 MJ
Radius 0.963 R 0.1737 RJ
Relaxation factor (s-1) 5020050-200 $0.2-7\times 10^{-7}$
Moment of Inertia ($\times\ mR^{2}$) 0.07 0.33
@@
Rotation Period (days) 42.7 -
Semi-major axis (AU) 0.015439
Eccentricity 0.028